OpenStudy (anonymous):

what are the zeros of the function below:

OpenStudy (anonymous):

\[\frac{ 4x^2-36x }{ x-9 }\]

OpenStudy (anonymous):

@ranga

OpenStudy (radar):

x=0 would result in a 0, of course if x=9 there would be discontinuity.

OpenStudy (anonymous):

so the answer choices are : x = -9 x = 0 x= 9 x = 0 and x = 9

OpenStudy (anonymous):

is it x = 9?

OpenStudy (anonymous):

lol just notices i posted this is OS feedback instead of math -_-

OpenStudy (radar):

No if x=9, it would result in a division by 0 and be undefined. x=0 would result in a 0.

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

x = 0?

OpenStudy (ranga):

x = 9 is NOT in the domain of the given function. The function is not defined at x = 9. So the solution should not include x = 9.

OpenStudy (anonymous):

so its a right, because thats the only solution that works

OpenStudy (ranga):

x = 0 is the solution. (choice B)

OpenStudy (anonymous):

oh . i see

OpenStudy (anonymous):

can you help me with 4 more?

OpenStudy (anonymous):

they have to do with discontinuities

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Post one more here and I will try.

OpenStudy (anonymous):

what is the discontinuity of the function \[f(x) = \frac{ x^2 - 3x -28 }{ x+4 }\]

OpenStudy (ranga):

If the denominator goes to 0 then the function is not defined because you cannot divide by zero. For what value of x does the denominator go to zero?

OpenStudy (anonymous):

hmm 0

OpenStudy (charlotte123):

Welcome To OpenStudy! @troll?

OpenStudy (anonymous):

im not new charlotte this is my alt acc since my main is buggy

OpenStudy (anonymous):

is it 0 range?

OpenStudy (charlotte123):

Ah I see LOL Well - Welcome Back!

OpenStudy (anonymous):

thanks

OpenStudy (ranga):

For what value of x does x+4 become zero?

OpenStudy (anonymous):

-4

OpenStudy (ranga):

Yes. x = -4 is the point of discontinuity. The function is not defined at x = -4.

OpenStudy (anonymous):

ok so what coordinate is the discontinuity ?

OpenStudy (anonymous):

(4, -3). is that it?

OpenStudy (ranga):

How did you get (4,-3)? What are the choices?

OpenStudy (anonymous):

a. (-7, -14) b. (7, 0) c. (4, -3) d. (-4, -11)

OpenStudy (ranga):

We just found at x = -4 there is a discontinuity. Which choice has -4 as the x coordinate?

OpenStudy (anonymous):

d.

OpenStudy (ranga):

Yes. D is the answer. But if you are interested in knowing how the y-coordinate is -11 we can go over that if necessary.

OpenStudy (anonymous):

sure

OpenStudy (ranga):

\[f(x) = \frac{ x^2-3x-28 }{ x+4 } = \frac{ (x-7)(x+4) }{ (x+4) } = x - 7\] The function is not defined at x = -4. But it is defined everywhere else. In calculus they will teach you limit. As x approaches -4, f(x) approaches -11 because you can put x = -4 in the simplified expr3ession above where f(x) = x - 7 as x gets close to -4, f(x) will get close to (-4) - 7 = -11. So the y value is -11 at x = -4. (-4,-11) is the point of discontinuity.

OpenStudy (anonymous):

cool i think i grasp it now. so the x value is the one that is important to discontinuity?

OpenStudy (ranga):

Yes. That is where the function is undefined. The y value is not that important.

OpenStudy (anonymous):

ok so to find the discontinuity just find the x value that makes it equal 0?

OpenStudy (ranga):

One way a discontinuity occurs is division by zero. If a function has an expression in the denominator, find out for what value of x the denominator will go to 0. Those x values will be the points of discontinuity. Other ways discontinuities occur is if the function is undefined for certain values of x such as the square root function which is not defined for negative numbers. Or a logarithm function that is not defined for 0 or negative values, etc..

OpenStudy (anonymous):

woah thats complicated… lol . so for this expression the point of discontinuity is also -4, right? \[f (x) = \frac{ 2x^2+5x-12 }{ x+4 }\]

OpenStudy (ranga):

correct.

OpenStudy (anonymous):

but this question is asking for the zeros too. the choices are a. zero at (\[(\frac{ 3 }{ 2}, 0)\] b. zero at

OpenStudy (anonymous):

 for choice b